1. 3/8 divided by 2/9
2. 1/2 - 2/5
3. 1 divided by 1/3
4. 1/5 + 1/6
5. What is 8% of 30?
I would greatly appreciate it if you could help me out with step by step directions i am visual learner and that's how I learn best. Thank you hope to hear from you soon.

Answer by rfer(16322)

1)
3/8*9/2=27/16=1 11/16
2)
5/10-4/10=1/10
3)
1*3/1=3/1=3
4)
6/30+5/30=11/30
5)
30*0.08=2.4